### v1

@@ 4.1 Mono Rule Miner
Early Termination (Observation #2): Our second observation arises from our definition of rule validity. Since a rule’s validity hinges on the validity of all ancestors, then if a rule candidate is found to be invalid, then none of its descendants can possibly be valid. This property is analogous to the Apriori property [3] from the data mining literature, and is key to “pruning” large subtrees of the lattice search space. Upon encountering an invalid rule, we can cache a record of its invalidity and avoid evaluating (i.e., performing model inference for) the child rules if and when we reach them during traversal. Furthermore, once a level has no valid rules, the entire search can terminate as all valid rules must have been found.
@@ 4.2 Dual Rule Miner
As the name implies, the limitation of the Mono Rule Miner is that only rules of a single type are mined in its sole pass. In settings where comprehensive tests are demanded (e.g., LLM red teaming), obtaining rules for a pair of reciprocal rules types enables deeper understanding of the RAG sources and their influence. With rules of two reciprocal types, a user can compare whether a set of sources $\mathbf{s^{\prime}}$ was sufficient, both when retained and omitted, to induce an LLM output that meets a single target condition. If $\mathbf{s^{\prime}}$ is implicated in both types of rules, then users can be more confident in its absolute importance. However if $\mathbf{s^{\prime}}$ is implicated in only one rule type, users can infer that its influence is relative to the presence (or absence) of the other sources.
To find rules of both types using Mono, each would require a separate pass, and may utilize different output predicates. Despite potentially different predicates, the underlying BFS still navigates the same lattice, in the same order, for the same input set $\mathbf{s}$, regardless of the chosen type. Algorithmically, the key difference between the two types stems from the interpretation of the “current subset;” i.e., how it will be applied. For the current subset $\mathbf{s^{\prime}}\in\mathbf{s}$, a retention rule must retain $\mathbf{s^{\prime}}$, but an omission rule must omit $\mathbf{s^{\prime}}$.
@@ 5 Experimental Evaluation
In this section, we conduct an experimental evaluation to analyze the quality of our rule mining algorithms. We begin with a qualitative evaluation that describes a RAG case study in the domain of healthcare. Then, we present a quantitative evaluation that studies our rule miners applied to a real-world RAG QA problem. We analyze the efficiency improvements gained by using our algorithms. Finally, in an effort to contrast these findings with theoretical boundaries, we then conduct a similar evaluation with synthetic data.
5.1 RAG Case Study
Figure 3: A lattice corresponding to the case study presented in Section 5.1. Here, the minimal valid rule (lowest valid rule in the lattice, as per Definition 8) states that “if D2 and D4 are retained, then the LLM produces a response that contains misinformation.”
@@ 5.1 RAG Case Study
The user is investigating a concerning trend that several doctors have identified: when posed with questions related to Long COVID, the chatbot often returns a misinformed answer that downplays the efficacy of pulmonary rehabilitation, despite clinical consensus of its benefits. To dig deeper, the technical user wishes to understand how the LLM’s answer changes when seeded with different documents provided via RAG. Specifically, they wonder whether certain “bad” documents lead the LLM to produce this misinformed answer, and therefore seek rules that link specific documents to this line of misinformation.
@@
The user decides to debug a specific question, “Is pulmonary rehabilitation effective for treating Long COVID symptoms?,” focusing first on the documents returned by the system. Upon review, the user confirms that the retriever is functioning correctly, and that the retrieved documents are relevant to the query. The user suspects that, despite their perceived relevance to the question, one or more of the retrieved documents may contain misinformation, which may be repeated by the LLM in turn. Therefore, the user seeks rules that will expose problematic documents and link them to misinformed answers.
To generate the rules, the user defines an output predicate that uses a different LLM to judge whether the response contains known medical misinformation. The user employs Mono Rule Miner to find all valid retention rules. As illustrated in Figure 3, several rules are returned, all originating from a single minimal rule that states the following: whenever documents $\left\{\text{D2},\text{D4}\right\}$ are retained, the LLM response will always contain misinformation. This suggests that, although neither D2 nor D4 are sufficient individually to induce a misinformed answer (i.e., there is no valid rule for $\left\{\text{D2}\right\}$ or $\left\{\text{D4}\right\}$), their combined presence can convince the LLM of its legitimacy.
With these new insights, the technical user can now flag these two documents for further review by a medical expert, and temporarily remove it from the database. After the expert revises the document to remove the problematic content, the same rule-based explanations can aid in validating a successful redeployment. By running the rule miner again, the technical user should find that no rules implicating misinformed responses exist, ensuring that this misinformation is less likely to propagate to end users.
5.2 Quantitative Evaluation: LLM Question Answering
@@ 5.2.1 Experimental Setup
To test our rule miners in a modern setting, we design an experiment that tasks a RAG system with question answering (QA), seeking rules to explain its answer to specific questions. We focus specifically on a scenario where a factual question is posed to an LLM, alongside a set of knowledge sources for reference (i.e., to help answer the question), and is prompted to return the correct answer. In an attempt to rule out the influence of pre-trained LLM knowledge, our prompt instructs the LLM to answer the question using only the information from these sources. For this evaluation, we use OpenAI’s gpt-4o-mini-2024-07-18 LLM as our model $M$. To realize this experimental setting, we utilize the HotpotQA dataset [39], a QA benchmark for LLMs and other QA systems.
@@
HotpotQA tests a model’s ability to answer multi-hop questions, which require consideration of multiple independent pieces of information (“hops”) to answer. For example, a multi-hop question might ask, “What is the chemical element named after the physicist who developed the theory of relativity?” Logically, answering this question requires one to first learn that Albert Einstein primarily developed the theory of relativity, then learn that the element named after him is Einsteinium. Each HotpotQA question is labeled with a single time-invariant correct answer, a set of “context” sources (i.e., potentially relevant documents), and a list of “supporting facts.” The supporting facts are pointers to specific sentences in the context sources, and collectively represent a minimal set of facts (sentences) necessary to derive the correct answer. Naturally, each question has at least two facts.
@@
As formalized in Section 3.2, our rule miners require a definition for the input set $\mathbf{s}$ and the output predicate(s) $O$. To decide on a formulation, we must motivate our search for rules. We adopt an intuitive problem definition wherein a user wishes to understand how the context sources provided to the LLM (via HotpotQA) affect the correctness of the LLM’s answer.
Therefore, when posed with a specific HotpotQA question, we define $\mathbf{s}$ as the set of sources that we provide to the LLM. The question and other input components are defined under $\mathbf{c}$, since they are not the focus of our present formulation. To maintain consistent granularity across context sources and supporting facts, our definition of $\mathbf{s}$ treats each sentence from each context source as a unique source. When these sentences (sources) are inserted into a RAG prompt, we ensure that they appear in their original order (i.e., we partially reconstruct the context “document” sources from the sentences present). We provide different numbers of sources throughout the experiment, typically ensuring that some number of supporting sources are included.
@@
Meanwhile, we define two complementary output predicates, $O_{correct}$ and $O_{incorrect}$, as described in Section 3.1. The predicate function first attempts to judge a predicted answer’s correctness via simple regular expression matching against the ground truth answer. If there is no match, it asks an LLM (also gpt-4o-mini-2024-07-18) to judge correctness. Under this formulation, our algorithms will mine the following two types of rules:
@@
    Retention rules: Identify source subsets that, when retained within a RAG prompt, always induce a correct answer from the LLM.
@@
    Omission rules: Identify source subsets that, when omitted from a RAG prompt, always induce an incorrect answer from the LLM.
The goal of this quantitative evaluation is to measure the average improvement of our rule miners, which leverage novel pruning and early termination strategies, as compared to a naive method that does not. To quantify the degree of pruning, it suffices to measure the average proportion of the lattice nodes that an algorithm actually visits. In this setting, visiting a node implies executing a single LLM inference call. Expressing this as a proportion allows us to average and consolidate results over lattices of varying sizes.
We hypothesize that the effectiveness of our pruning depends on the lattice size. In smaller lattices, there are fewer nodes to prune, so the amount of nodes visited is similar to that of a naive algorithm. However in large lattices, each node pruned invalidates a larger subtree in the lattice, leading to greater overall pruning. Intuitively, more nodes will exist in a larger lattice, meaning that more nodes (by count) could be pruned, but does the proportion of nodes eligible for pruning also increase? This is the hypothesis that we test in this experiment.
To test this hypothesis, we proceed as follows. For each mining algorithm under test (Mono and Dual), we measure the proportion of nodes pruned over increasing numbers of sources in $\mathbf{s}$ (i.e., increasing lattice sizes), averaged over 120 HotpotQA examples (questions). One potential confounder is the true number of supporting facts required to piece together the correct answer. We hypothesize also that this number affects the number of valid rules in our RAG setting, and that in turn, it affects the number of nodes that can be pruned. To verify this hypothesis, we pose only those HotpotQA questions that have exactly three supporting facts, then test input source sets $\mathbf{s}$ where $0\leq|\mathbf{s}|\leq 6$. When $|\mathbf{s}|\leq 3$, we ensure that $\mathbf{s}$ consists only of supporting facts. Thus, our results will illustrate the pruning improvements when the LLM has too few supporting sources, all supporting sources, and extra sources in addition to all supporting.
@@ 5.2.2 Experimental Results: Pruning Improvements
As illustrated in Figure 4(a), the results support our hypothesis. For both Mono and Dual rule miners, the degree of pruning (i.e., the inverse of “Average Proportion of Lattice Explored”) is positively correlated with increasing lattice size (“Number of Sources”). As hypothesized, smaller lattices (i.e., few supporting sources) offer less opportunity for pruning; however, Mono prunes at an increasing rate as more supporting sources are included. We observe that Mono prunes more retention rule candidates than omission candidates for small lattice sizes; however, both progressively converge at the point where $\mathbf{s}$ includes all three supporting sources. After this point (when more sources are added), the proportion of pruning continues to increase at a slowing rate.

### v2

@@ 4.1 Mono Rule Miner
Early Termination (Observation #2): Our second observation arises from our definition of rule validity. Since a rule’s validity hinges on the validity of all ancestors, then if a rule candidate is found to be invalid, then none of its descendants can possibly be valid. This property is analogous to the downward closure property (Apriori property) from the data mining literature [apriori], and is key to “pruning” large subtrees of the lattice search space. Upon encountering an invalid rule, we can cache a record of its invalidity and avoid evaluating (i.e., performing model inference for) the child rules if and when we reach them during traversal. Furthermore, once a level has no valid rules, the entire search can terminate as all valid rules must have been found.
@@ 4.2 Dual Rule Miner
As the name implies, the limitation of the Mono Rule Miner is that only rules of a single type are mined in its sole pass. In settings where comprehensive tests are demanded (e.g., LLM red teaming), obtaining rules for a pair of reciprocal rule types enables deeper understanding of the RAG sources and their influence. With rules of two reciprocal types, a user can compare whether a set of sources $\mathbf{s^{\prime}}$ was sufficient, both when retained and omitted, to induce an LLM output that meets a single target condition. If $\mathbf{s^{\prime}}$ is implicated in both types of rules, then users can be more confident in its absolute importance. However if $\mathbf{s^{\prime}}$ is implicated in only one rule type, users can infer that its influence is relative to the presence (or absence) of the other sources.
To find rules of both types using Mono, each would require a separate pass, and may utilize different output predicates. Despite potentially different predicates, the underlying BFS still navigates the same lattice, in the same order, for the same input set $\mathbf{s}$, regardless of the chosen type. Algorithmically, the key difference between the two types stems from the interpretation of the “current subset;” i.e., how it will be applied. For the current subset $\mathbf{s^{\prime}}\subseteq\mathbf{s}$, a retention rule must retain $\mathbf{s^{\prime}}$, but an omission rule must omit $\mathbf{s^{\prime}}$.
@@ 5 Experimental Evaluation
In this section, we conduct an experimental evaluation [^2] The code to reproduce both quantitative experiments is available at https://github.com/joelrorseth/rag-llm-rule-miner. to analyze the efficiency and quality of our rule mining algorithms. We begin with a qualitative evaluation that describes a RAG case study in the domain of healthcare. Then, we present a quantitative evaluation studying the efficiency and scalability of our rule miners, applied in both real-world and synthetic settings. This is followed by another case study demonstrating a large-scale application of our rule miners in the domain of AI safety. Lastly, we present another quantitative evaluation studying the robustness of the rule miners and the rules they find across different commercial LLMs.
5.1 RAG Healthcare Case Study
Figure 3: A lattice corresponding to the case study presented in Section 5.1. Here, the minimal valid rule (lowest valid rule in the lattice, as per Definition 8) states that “if D2 and D4 are retained, then the LLM produces a response that is inconsistent with clinical evidence.”
@@ 5.1 RAG Healthcare Case Study
The user is investigating a concerning trend that several doctors have identified: when posed with questions related to Long COVID, the chatbot often returns a misinformed answer that downplays the efficacy of pulmonary rehabilitation, despite clinical consensus of its benefits. The user suspects that one or more documents being retrieved and provided to the LLM contain medical misinformation, and that the LLM is ultimately grounding its responses in this misinformation. To dig deeper, the technical user wishes to understand how the LLM’s answer changes when seeded with different documents provided via RAG. Specifically, they wonder whether certain document combinations lead the LLM to produce responses inconsistent with the authoritative clinical evidence, and therefore seek rules that link specific documents to this inconsistency.
@@
The user decides to debug a specific question, “Is pulmonary rehabilitation effective for treating Long COVID symptoms?,” focusing first on the documents returned by the system. Upon review, the user confirms that the retriever is functioning correctly, and that the retrieved documents are relevant to the query. The user suspects that, despite their relevance, one or more of the retrieved documents may contain conflicting or inaccurate information that causes the LLM to produce a response inconsistent with clinical evidence. Therefore, the user seeks rules that will identify these problematic documents and link them to the inconsistent answers.
To generate the rules, the user defines an output predicate that uses a secondary LLM to judge the inconsistency of the response w.r.t. authoritative clinical evidence (supplied by the user). The user employs Mono Rule Miner to find all valid retention rules. As illustrated in Figure 3, several rules are returned, all originating from a single minimal rule that states the following: whenever documents $\left\{\text{D2},\text{D4}\right\}$ are retained, the LLM response is inconsistent. This suggests that, although neither D2 nor D4 are sufficient individually to induce an inconsistent answer (i.e., there is no valid rule for $\left\{\text{D2}\right\}$ or $\left\{\text{D4}\right\}$), their combined presence is.
With these new insights, the technical user can now flag these two documents for further review by a medical expert, and temporarily remove one or both from the database to prevent inconsistent answers. After the expert revises the documents to ensure the content is consistent with authoritative clinical evidence, the same rule-based explanations can aid in validating a successful redeployment. By running the rule miner again, the technical user should find that no rules implicating inconsistent answers exist, ensuring that the system is more likely to remain faithful to the clinical consensus.
5.2 Quantitative Evaluation: Rule Miner Efficiency
In this section, we present a quantitative experiment that evaluates the pruning improvements and scalability of our rule miners, measured over both real-world and synthetic data. We analyze rule miners’ efficiency both absolutely and relatively, discussing practical implications and directions for future improvement.
@@ 5.2.1 Experimental Setup
To test our rule miners in a modern setting, we design an experiment that tasks a RAG system with question answering (QA), seeking rules to explain its answer to specific questions. We focus specifically on a RAG QA scenario where a factual question is posed to an LLM along with a set of knowledge sources for reference. The LLM, $M$, is prompted to return the correct answer if it can be inferred from the knowledge sources, or “N/A” otherwise. The full implementation details of this “LLM-under-test” are provided in Appendix A. To realize this experimental setting, we utilize the HotpotQA dataset [hotpotqa], a QA benchmark for LLMs and other QA systems.
@@
HotpotQA tests a model’s ability to answer multi-hop questions, which require consideration of multiple independent pieces of information (“hops”) to answer. HotpotQA questions have exactly two hops. An example multi-hop question with two hops might ask, “What is the chemical element named after the physicist who developed the theory of relativity?” Logically, answering this question requires one to first learn that Albert Einstein primarily developed the theory of relativity, then learn that the element named after him is Einsteinium. Each HotpotQA question is labeled with a single ground truth answer, a set of necessary sources required to infer the answer (constructed from HotpotQA “supporting facts”), and a set of distractor sources that are inconsequential to the determination of the answer. The sets of necessary and distractor sources are disjoint, and there are exactly $k=2$ necessary sources for each question (one for each hop).
@@
As formalized in Section 3.2, our rule miners require a definition for the input set $\mathbf{s}$ and the output predicate(s) $O$. To decide on a formulation, we must motivate our search for rules. We adopt an intuitive problem definition wherein a user wishes to understand how the context sources provided to the LLM (via HotpotQA) affect the correctness of the LLM’s answer. Therefore, when posed with a specific HotpotQA question, we define $\mathbf{s}$ as the set of sources that we provide to the LLM. The question and other input components constitute $\mathbf{c}$, since they are not the focus of our present formulation.
@@
Meanwhile, we define two complementary output predicates, $O_{consistent}$ and $O_{inconsistent}$, which attempt to classify answers predicted by the LLM based on their consistency w.r.t. the ground truth answer. The predicate function first attempts to judge a predicted answer’s correctness via deterministic string matching against the ground truth answer. If the answers do not match, they are posed to an LLM (gpt-5-mini-2025-08-07) that judges the equivalency of the candidate answer w.r.t. the ground truth answer. The full implementation details for the string matching and “LLM-as-a-judge” procedures are provided in Appendix A. Under this formulation, our algorithms will mine the following two types of rules:
@@
    Retention rules: Identify source subsets that, when retained within a RAG prompt, always induce an answer consistent with the ground truth answer.
@@
    Omission rules: Identify source subsets that, when omitted from a RAG prompt, always induce an answer inconsistent with the ground truth answer.
The goal of this quantitative evaluation is to measure the efficiency of our rule miners, which leverage novel pruning and early termination strategies, as compared to a naive method that does not. We propose that the efficiency of a rule miner can be quantified by the average proportion of lattice nodes that the rule miner actually visits. In this setting, “visiting” a node implies executing a single LLM inference call, therefore the number of calls executed by the rule miner can also quantify its efficiency. Expressing the number of nodes visited as a proportion allows these measurements to be averaged and consolidated over lattices of varying sizes (i.e., varying numbers of possible nodes).
Using these two metrics, we proceed as follows. We draw 50 QA pairs at random from the HotpotQA training set (hotpot_train_v1.1.json), which will ultimately be posed to a single “LLM-under-test” $M$ (gpt-5-mini-2025-08-07). For each question, we task each of our rule miners (Mono seeking retention rules, Mono seeking omission rules, and Dual seeking both) to mine all rules arising from posing said question to $M$. Given that $k=2$ (i.e., only two sources should be consequential to consistency), we anticipate more opportunity for pruning as additional distractor sources are considered and $|\mathbf{s}|$ increases. Therefore, we pose each question repeatedly to $M$ with increasingly large $\mathbf{s}$ (spanning $1\leq|\mathbf{s}|\leq 10$), exhausting necessary sources before drawing additional distractor sources.
To contrast the results of the real-world HotpotQA setting, we also compute results in a synthetic setting. In this setting, we control for several real-world factors that affect the true validity of rules found. We create a dataset containing mock sources (both necessary and distractor), then instantiate a mock LLM-as-a-judge and LLM-under-test $M$ that use them to predict consistent vs. inconsistent perfectly. In effect, this controls for the non-determinism and stochasticity of both LLMs, as well as for their propensity to make inaccurate and inconsistent predictions. We aggregate results over 1000 synthesized examples, and otherwise perform the same steps as in the HotpotQA setting. In the following, we discuss results and contrast them in the HotpotQA vs. synthetic settings.
@@ 5.2.2 Experimental Results: Pruning Improvements
(a) Rule miner lattice exploration as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Rule miner lattice exploration as evaluated in a synthetic setting that controls for stochasticity and consistency.